OpenStudy (anonymous):

Find the perihelion and aphelion distances of Saturn given it travels in an elliptical orbit with the Sun at one focal point and has a semi-major axis of 1427 Gm and an eccentricity of .056

OpenStudy (anonymous):

eccentricity = c/a aphelion= a+c perihelion = a-c But I don't know what the semi-major axis is

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

the letter for an ellipse relate to a rt triangle; but they are mixed up from the usual pythag thrm set up

OpenStudy (amistre64):



OpenStudy (anonymous):

Which letter is the semi-major axis?

OpenStudy (amistre64):

a is the semi (half-part) of the major (biggest) axis

OpenStudy (anonymous):

So a, right?

OpenStudy (amistre64):

b is the minor; and c is the focal distance from center

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

your welcome, and good luck